# Distance Between Stochastic Linear Systems

### v1

@@ 3.3 Type-$q$ Distance Between Systems $P_{1}$ and $P_{2}$
To address the above shortcoming, we propose the type-$q$ Wasserstein distance between systems by taking into account their distributions $\mathbb{P}_{P_{1}(\omega)}$ and $\mathbb{P}_{P_{2}(\omega)}$ for every frequency $\omega\in\Omega$.
@@
$\mathrm{d}_{q}\left(P_{1},P_{2}\right)$ $:=\sup_{\omega\in\Omega}W^{q}_{q}\left(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\right)$
(26) $=\sup_{\omega\in\Omega}\,\underset{\pi_{\omega}\in\Pi\left(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\right)}{\inf}\int_{\mathcal{R}_{P_{1}}(\omega)\times\mathcal{R}_{P_{2}}(\omega)}\mathrm{d_{chord}}(r_{1},r_{2})^{q}\,\pi_{\omega}(dr_{1},dr_{2}).$
@@ 3.3.1 Upper Bound on $\mathrm{d}_{q}\left(P_{1},P_{2}\right)$
(27) $\mathrm{d}_{q}(P_{1},P_{2})\leq\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})^{q}.$
Proof 3.10.
Given $q\geq 1$, fix any $\omega\in\Omega$. Let $\mu:=\mathcal{U}(\mathcal{R}_{P_{1}}(\omega))$ and $\nu:=\mathcal{U}(\mathcal{R}_{P_{2}}(\omega))$ be the uniform distributions over the respective compact support sets. Then, the chordal distance metric induced type-$q$ Wasserstein distance using (3.6) is given by
@@
Taking supremum over all $\omega\in\Omega$, we get (27) and the proof is complete.
@@ 3.3.2 Lower Bound on $\mathrm{d}_{q}\left(P_{1},P_{2}\right)$
Having obtained an upper bound for the distance metric $\mathrm{d}_{q}(P_{1},P_{2})$ in Theorem˜3.9, we now proceed below to get a lower bound using triangle inequality based arguments. Specifically, we will leverage the nominal distance and the expected deviation of the random plant instances of each systems from their respective nominal models to arrive at a lower bound for the proposed distance measure.
@@
$\mathrm{d}_{q}(P_{1},P_{2})$
(28) $\geq\sup_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})}\left(\mathrm{d_{chord}}(\overline{R}_{1}(\omega),\overline{R}_{2}(\omega))-\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\Delta_{\mathrm{nom}}(\omega)\right]\right)^{q}_{+},$
@@
(29) $\Delta_{\mathrm{nom}}(\omega):=\sum^{2}_{\ell=1}\mathrm{d_{chord}}(r_{\ell}(\omega),\overline{R}_{\ell}(\omega)).$
Proof 3.12.
Given $q\geq 1$, fix any frequency $\omega\in\Omega$ and let $\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})$ be any admissible joint distribution at that frequency. For the brevity of notation, we will drop the $\omega$ argument from the quantities of interests when it is obvious that quantities are function of frequency $\omega$. For any $r_{1}\in\mathcal{R}_{P_{1}}$ and $r_{2}\in\mathcal{R}_{P_{2}}$, we can define $\Delta_{\mathrm{nom}}$ that denotes the sum of deviation (measured in terms of the chordal distance metric) of any perturbed models from their respective nominal models using (29) and it is a random variable. Using the triangle inequality for the chordal metric $\mathrm{d_{chord}}$ using $\Delta_{\mathrm{nom}}$, we see that
@@
$\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\mathrm{d_{chord}}(r_{1},r_{2})^{q}\right]$ $\geq\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\Delta_{\mathrm{nom}}\right)^{q}_{+}\right]$
@@
where we applied Jensen’s inequality (for convex function $f(x)=(x)^{q}_{+}$ for $q\geq 1$, Jensen’s inequality implies $\mathbb{E}[(x)^{q}_{+}]\geq(\mathbb{E}[x])^{q}_{+}$) on the right hand side. Note that $\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\mathrm{d_{chord}}(r_{1},r_{2})^{q}\right]$ is simply equal to $\int_{\mathcal{R}_{P_{1}}\times\mathcal{R}_{P_{2}}}\mathrm{d_{chord}}(r_{1},r_{2})^{q}\,d\pi_{\omega}$. Taking the infimum over all joint distributions $\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})$ on both sides, we get
$W_{q}^{q}(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})\geq\inf_{\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})}\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\Delta_{\mathrm{nom}}\right]\right)^{q}_{+}.$
Taking supremum over all $\omega\in\Omega$ on both sides yields (3.11).
@@
We believe that a similar lower bound like the one in (3.11) for the $\mathrm{d}_{q}(P_{1},P_{2})$ can be obtained using [21, Theorem 14.60] by adapting the integral-infimum interchange theorem to supremum, where interchange of supremum and integration happens under conditions favouring Fubini-type arguments. Similarly, we can use Dobrushin inequality given any feasible candidate transport plan to obtain a simple yet conservative upper bound for the distance measure. We leave both the expositions as future works.
@@ 3.4 $\mathbb{P}_{P_{1}(\omega)},\mathbb{P}_{P_{2}(\omega)}$ Being Empirical Distributions
(30) $\mathbb{P}_{P_{\ell}(\omega)}$ $=\frac{1}{N}\sum^{N}_{i=1}\boldsymbol{\delta}_{\hat{P}^{(i)}_{\ell}(j\omega)},\quad\ell=1,2,$
@@
$\mathrm{d}_{q}(P_{1},P_{2})$ $=\sup_{\omega\in\Omega}\,W^{q}_{q}\left(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\right)$
(31) $=\sup_{\omega\in\Omega}\,\inf_{\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})}\sum^{N}_{i=1}\sum^{N}_{k=1}\mathrm{d_{chord}}\left(\hat{R}^{(i)}_{1}(j\omega),\hat{R}^{(k)}_{2}(j\omega)\right)^{q}\,\pi_{\omega}(i,k),$
@@
Proof 3.15.
Adapting the distance calculation from (3.6) to the Wasserstein distance between empirical distributions $\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}$ using the chordal metric given by (22) yields the result (3.14).
The continuity in the frequency $\omega\in\Omega$ dimension still makes the exact computation of (3.14) hard. One way to approximately address this problem is to discretize the frequency dimension as finely as possible and deploy numerical approximation methods. That is, we can define the discretized frequency space with $M\in\mathbb{N}$ points as $\Omega_{M}=\left\{\omega_{i}\in\Omega\right\}^{M}_{i=1}$. Then, the distance in (3.14) can be approximated as
(32) $\hat{\mathrm{d}}_{q}(P_{1},P_{2})$ $=\max_{\omega\in\Omega_{M}}\,\inf_{\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})}\sum^{N}_{i=1}\sum^{N}_{k=1}\mathrm{d_{chord}}(\hat{R}^{(i)}_{1}(j\omega),\hat{R}^{(k)}_{2}(j\omega))^{q}\,\pi_{\omega}(i,k).$
@@
(33) $\mathbb{P}_{R_{\ell}(\omega)}=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\delta}_{R_{\ell}^{(i)}(j\omega)},\quad\ell=1,2.$
@@
(34a) $C(\omega)$ $:=\sup_{i\in[1,N],k\in[1,N]}\mathrm{d_{chord}}\left(R_{1}^{(i)}(j\omega),R_{2}^{(k)}(j\omega)\right).$
@@
(35) $\hat{d}_{q}(P_{1},P_{2})\leq\max_{\omega\in\Omega_{M}}(C(\omega))^{q}.$
Proof 3.17.
@@
$W_{q}^{q}(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})\leq\sum_{i=1}^{N}\sum_{k=1}^{N}(C(\omega))^{q}\,\hat{\pi}_{\omega}(i,k)=(C(\omega))^{q}\sum_{i=1}^{N}\sum_{k=1}^{N}\,\hat{\pi}_{\omega}(i,k)=(C(\omega))^{q}.$
Taking infimum over the set of all possible joint distributions $\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})$ and subsequently taking maximum over $\omega\in\Omega_{M}$ on both sides gives (35).
Following the similar thinking of Theorem˜3.11, we will now obtain a lower bound for the empirical version of the distance measure $\hat{d}_{q}(P_{1},P_{2})$ based on the triangle inequality involving the deviation of the random models of each system from their nominal models.
@@
(36) $\Delta^{(ik)}_{\mathrm{nom}}(\omega):=\mathrm{d_{chord}}\left(\omega;R_{1}^{(i)},\overline{R}_{1}\right)+\mathrm{d_{chord}}\left(\omega;R_{2}^{(k)},\overline{R}_{2}\right),\quad\text{and},$
(37) $\hat{d}_{q}(P_{1},P_{2})\geq\max_{\omega\in\Omega_{M}}\inf_{\pi_{\omega}\in\Pi(\omega;\mathbb{P}_{R_{1}},\mathbb{P}_{R_{2}})}\left(\mathrm{d_{chord}}(\omega;\overline{R}_{1},\overline{R}_{2})-\sum^{N}_{i=1}\sum^{N}_{k=1}\Delta^{(ik)}_{\mathrm{nom}}(\omega)\,\pi_{\omega}(i,k)\right)^{q}_{+}.$
Proof 3.19.
We will follow the arguments in proof of Theorem˜3.11. Let us fix a frequency $\omega\in\Omega_{M}$. As usual, for the brevity of notation, we will drop the $\omega$ argument when things are obvious. Using the triangle inequality for the chordal distance metric $\mathrm{d_{chord}}$ on the Riemann sphere involving the deviation of system models from their respective nominal models, we see that for every sample $R_{1}^{(i)}\in\mathrm{supp}(\mathbb{P}_{R_{1}})$ and $R_{2}^{(k)}\in\mathrm{supp}(\mathbb{P}_{R_{2}})$ with $i,k=\{1,\dots,N\}$, we see that
(38) $\mathrm{d_{chord}}\left(R_{1}^{(i)},R_{2}^{(k)}\right)\geq\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\Delta^{(ik)}_{\mathrm{nom}}\right)_{+}$,
@@
$\sum^{N}_{i=1}\sum^{N}_{k=1}\mathrm{d_{chord}}\left(R_{1}^{(i)},R_{2}^{(k)}\right)^{q}\,\pi_{\omega}(i,k)$ $\geq\sum^{N}_{i=1}\sum^{N}_{k=1}\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\Delta^{(ik)}_{\mathrm{nom}}\right)^{q}_{+}\,\pi_{\omega}(i,k)$
$\geq\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\sum^{N}_{i=1}\sum^{N}_{k=1}\Delta^{(ik)}_{\mathrm{nom}}\pi_{\omega}(i,k)\right)^{q}_{+},$
where we again used the Jensen’s inequality as we did before in the proof of Theorem˜3.11. Subsequently, taking the infimum over all possible joint distributions $\pi_{\omega}\in\Pi(\omega;\mathbb{P}_{R_{1}},\mathbb{P}_{R_{2}})$, and maximizing over $\omega\in\Omega_{M}$ on both sides yields (37).

### v2

@@ 3.3 Type-$q$ Distance Between Systems $P_{1}$ and $P_{2}$
We address the above shortcoming by proposing the type-$q$ Wasserstein distance between systems by taking into account their distributions $\mathbb{P}_{P_{1}(\omega)}$, $\mathbb{P}_{P_{2}(\omega)}$ for every frequency $\omega\in\Omega$. We denote the corresponding set of all possible joint distribution by $\Pi_{\omega}:=\Pi\left(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\right)$.
@@
$\mathrm{d}_{q}\left(P_{1},P_{2}\right)$
$:=\sup_{\omega\in\Omega}W^{q}_{q}\left(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\right)$
$=\sup_{\omega\in\Omega}\,\underset{\pi_{\omega}\in\Pi_{\omega}}{\inf}\int_{\mathcal{R}_{P_{1}}(\omega)\times\mathcal{R}_{P_{2}}(\omega)}\mathrm{d_{chord}}(r_{1},r_{2})^{q}\,\pi_{\omega}(dr_{1},dr_{2}).$ (26)
@@ 3.3.1 Upper Bound on $\mathrm{d}_{q}\left(P_{1},P_{2}\right)$
$\mathrm{d}_{q}(P_{1},P_{2})\leq\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})^{q}.$ (27)
Proof.
Given $q\geq 1$, fix any $\omega\in\Omega$. Let $\mu:=\mathcal{U}(\mathcal{R}_{P_{1}}(\omega))$ and $\nu:=\mathcal{U}(\mathcal{R}_{P_{2}}(\omega))$ be the uniform distributions over the respective compact support sets. Then, the chordal metric induced type-$q$ Wasserstein distance using (3.3) is given by
@@
Taking supremum over all $\omega\in\Omega$, we get (27) and the proof is complete. ∎
@@ 3.3.2 Lower Bound on $\mathrm{d}_{q}\left(P_{1},P_{2}\right)$
Having obtained an upper bound for the distance metric $\mathrm{d}_{q}(P_{1},P_{2})$ in Theorem 3.2, we now proceed below to get a lower bound using triangle inequality based arguments. We will leverage the nominal distance and the expected deviation of the random plant instances of each systems from their respective nominal models to arrive at a lower bound for the proposed distance measure.
@@
$\mathrm{d}_{q}(P_{1},P_{2})$ $\geq\sup_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi_{\omega}}\Delta_{\mathrm{dev}}(\omega),$ (28)
@@
$\Delta_{\mathrm{dev}}(\omega)$ $=\left(\mathrm{d_{chord}}(\overline{R}_{1}(\omega),\overline{R}_{2}(\omega))\!-\!\mathbb{E}\left[\Delta_{\mathrm{nom}}(\omega)\right]\right)^{q}_{+}$ (29a)
$\Delta_{\mathrm{nom}}(\omega)$ $=\sum^{2}_{\ell=1}\mathrm{d_{chord}}(r_{\ell}(\omega),\overline{R}_{\ell}(\omega)).$ (29b)
Proof.
Given $q\geq 1$, fix any frequency $\omega\in\Omega$ and let $\pi_{\omega}\in\Pi_{\omega}$ be any admissible joint distribution at that frequency. For the brevity of notation, we will drop the $\omega$ argument from the quantities of interests when it is obvious that quantities are function of frequency $\omega$. For any $r_{1}\in\mathcal{R}_{P_{1}}$ and $r_{2}\in\mathcal{R}_{P_{2}}$, we can define $\Delta_{\mathrm{nom}}$ that denotes the sum of deviation (measured in terms of the chordal distance metric) of any perturbed models from their respective nominal models using (29b) and it is a random variable. Using the triangle inequality for the $\mathrm{d_{chord}}$ using $\Delta_{\mathrm{nom}}$, we see that
@@
$\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\mathrm{d_{chord}}(r_{1},r_{2})^{q}\right]$
$\geq$ $\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\Delta_{\mathrm{nom}}\right)^{q}_{+}\right]$
@@
where we applied Jensen’s inequality (for convex function $f(x)=(x)^{q}_{+}$ for $q\geq 1$, Jensen’s inequality implies $\mathbb{E}[(x)^{q}_{+}]\geq(\mathbb{E}[x])^{q}_{+}$) on the right hand side. Note that $\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\mathrm{d_{chord}}(r_{1},r_{2})^{q}\right]$ is simply equal to $\int_{\mathcal{R}_{P_{1}}\times\mathcal{R}_{P_{2}}}\mathrm{d_{chord}}(r_{1},r_{2})^{q}\,d\pi_{\omega}$. Taking infimum over all joint distributions $\pi_{\omega}\in\Pi(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})$ on both sides, we get
$W_{q}^{q}(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})$
$\geq\inf_{\pi_{\omega}\in\Pi_{\omega}}\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})\!-\!\mathbb{E}_{(r_{1},r_{2})\sim\pi_{\omega}}\left[\Delta_{\mathrm{nom}}\right]\right)^{q}_{+}.$
Taking supremum over all $\omega\in\Omega$ on both sides yields (28) and the proof is complete. ∎
@@
We believe that a similar lower bound like the one in (28) for the $\mathrm{d}_{q}(P_{1},P_{2})$ can be obtained using [26, Theorem 14.60] by adapting the integral-infimum interchange theorem to supremum, where interchange of supremum and integration happens under conditions favouring Fubini-type arguments. Similarly, we can use Dobrushin inequality given any feasible candidate transport plan to obtain a simple yet conservative upper bound for the distance measure. We leave both the expositions as future works.
@@ 3.4 $\mathbb{P}_{P_{1}(\omega)},\mathbb{P}_{P_{2}(\omega)}$ Being Empirical Distributions
$\mathbb{P}_{P_{\ell}(\omega)}$ $=\frac{1}{N}\sum^{N}_{i=1}\boldsymbol{\delta}_{\hat{P}^{(i)}_{\ell}(j\omega)},\quad\ell=1,2,$ (30)
@@
$\footnotesize\mathrm{d}_{q}(P_{1},P_{2})\!=\!\sup_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi_{\omega}}\!\sum^{N}_{i=1}\sum^{N}_{k=1}\mathrm{d_{chord}}\left(\hat{R}^{(i)}_{1}(j\omega),\hat{R}^{(k)}_{2}(j\omega)\right)^{q}\!\!\pi_{\omega}(i,k),$ (31)
@@
Proof.
Adapting the distance calculation from (3.3) to the Wasserstein distance between empirical distributions $\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}$ using the chordal metric given by (22) yields the result (31). ∎
The continuity in the frequency $\omega\in\Omega$ dimension still makes the exact computation of (31) hard. One way to approximately address this problem is to discretize the frequency dimension as finely as possible and deploy numerical approximation methods. That is, we can define the discretized frequency space with $M\in\mathbb{N}$ points as $\Omega_{M}=\left\{\omega_{i}\in\Omega\right\}^{M}_{i=1}$. Then, the distance in (31) can be approximated as
$\footnotesize\hat{\mathrm{d}}_{q}(P_{1},P_{2})\!=\!\max_{\omega\in\Omega_{M}}\!\inf_{\pi_{\omega}\in\Pi_{\omega}}\!\sum^{N}_{i=1}\sum^{N}_{k=1}\mathrm{d_{chord}}(\hat{R}^{(i)}_{1}(j\omega),\hat{R}^{(k)}_{2}(j\omega))^{q}\pi_{\omega}(i,k).$ (32)
@@
$\mathbb{P}_{R_{\ell}(\omega)}=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\delta}_{R_{\ell}^{(i)}(j\omega)},\quad\ell=1,2.$ (33)
@@
$C(\omega)$ $:=\sup_{i\in[1,N],k\in[1,N]}\mathrm{d_{chord}}\left(R_{1}^{(i)}(j\omega),R_{2}^{(k)}(j\omega)\right).$ (34a)
@@
$\hat{d}_{q}(P_{1},P_{2})\leq\max_{\omega\in\Omega_{M}}(C(\omega))^{q}.$ (35)
Proof.
@@
$W_{q}^{q}(\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)})$ $\leq\sum_{i=1}^{N}\sum_{k=1}^{N}(C(\omega))^{q}\,\hat{\pi}_{\omega}(i,k)$
$=(C(\omega))^{q}\sum_{i=1}^{N}\sum_{k=1}^{N}\,\hat{\pi}_{\omega}(i,k)$
$=(C(\omega))^{q}.$
Taking infimum over the set of all possible joint distributions $\Pi_{\omega}$ and subsequently taking maximum over $\omega\in\Omega_{M}$ on both sides gives (35). ∎
Following the similar thinking of Theorem 3.3, we will now obtain a lower bound for the empirical version of the distance measure $\hat{d}_{q}(P_{1},P_{2})$ based on the triangle inequality involving the deviation of the random models of each system from their nominal models.
@@
$\Delta^{(ik)}_{\mathrm{nom}}(\omega)$ $:=\mathrm{d_{chord}}\left(\omega;R_{1}^{(i)},\overline{R}_{1}\right)+\mathrm{d_{chord}}\left(\omega;R_{2}^{(k)},\overline{R}_{2}\right),$ (36)
$\hat{d}_{q}(P_{1},P_{2})$ $\geq\max_{\omega\in\Omega_{M}}\inf_{\pi_{\omega}\in\Pi_{\omega}}\left(\Delta_{\mathrm{dev}}(\omega)\right)^{q}_{+},$ (37)
$\Delta_{\mathrm{dev}}(\omega)$ $=\mathrm{d_{chord}}(\omega;\overline{R}_{1},\overline{R}_{2})\!-\!\sum^{N}_{i=1}\sum^{N}_{k=1}\Delta^{(ik)}_{\mathrm{nom}}(\omega)\,\pi_{\omega}(i,k).$ (38)
Proof.
We will follow the arguments in proof of Theorem 3.3. Let us fix a frequency $\omega\in\Omega_{M}$. As usual, for the brevity of notation, we will drop the $\omega$ argument when things are obvious. Using the triangle inequality for the chordal distance metric $\mathrm{d_{chord}}$ on the Riemann sphere involving the deviation of system models from their respective nominal models, we see that for every sample $R_{1}^{(i)}\in\mathrm{supp}(\mathbb{P}_{R_{1}})$ and $R_{2}^{(k)}\in\mathrm{supp}(\mathbb{P}_{R_{2}})$ with $i,k=\{1,\dots,N\}$, we see that
$\mathrm{d_{chord}}\left(R_{1}^{(i)},R_{2}^{(k)}\right)\geq\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\Delta^{(ik)}_{\mathrm{nom}}\right)_{+}$, (39)
@@
$\sum^{N}_{i=1}\sum^{N}_{k=1}\mathrm{d_{chord}}\left(R_{1}^{(i)},R_{2}^{(k)}\right)^{q}\,\pi_{\omega}(i,k)$
$\geq\sum^{N}_{i=1}\sum^{N}_{k=1}\left(\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\Delta^{(ik)}_{\mathrm{nom}}\right)^{q}_{+}\,\pi_{\omega}(i,k)$
$\geq\left(\underbrace{\mathrm{d_{chord}}(\overline{R}_{1},\overline{R}_{2})-\sum^{N}_{i=1}\sum^{N}_{k=1}\Delta^{(ik)}_{\mathrm{nom}}\pi_{\omega}(i,k)}_{:=\Delta_{\mathrm{dev}}(\omega)}\right)^{q}_{+},$
where we used the Jensen’s inequality as we did before in the proof of Theorem 3.3. Subsequently, taking the infimum over all possible joint distributions $\pi_{\omega}\in\Pi(\omega;\mathbb{P}_{R_{1}},\mathbb{P}_{R_{2}})$, and maximizing over $\omega\in\Omega_{M}$ on both sides yields (37). ∎